# A Novel Density Profile for Isothermal Cores of Dark Matter Halos

### v3

@@ I Introduction
Following the initial thermalization phase and throughout the gravothermal collapse process, until reaching the short-mean-free-path regime when heat conduction is only effective at the surface of the core, the halo core exhibits self-similarity in both the density and velocity dispersion profiles. That is, throughout its evolution, the halo profiles can be characterized by specific values of scale density, radius, and velocity dispersion. This behavior has been well described theoretically [e.g., 29] and confirmed in simulations for both velocity-independent and velocity-dependent self-interacting cross sections [e.g., 31, 32]. Such self-similar nature provides the foundation for adopting particular functional forms to describe the density structure of halos. In this self-similar regime, the halo core develops characteristic plateaus in both the density and velocity dispersion profiles. The emergence of a central plateau in the density profile, commonly referred to as the “core configuration”, is a key feature of SIDM models and is instrumental in addressing the core-cusp problem, as discussed above. In this context, we also define the “isothermal-core configuration” as a state in which both the density and velocity dispersion remain approximately constant within the central region of the halo. Such behaviors of SIDM have been found in cosmological and isolated N-body simulations [e.g. 33, 22, 23, 24, 32]. In the outer regions of the halo, however, the number density and velocity dispersion of DM particles are significantly smaller than in the core, resulting in negligible scattering interactions over the age of the Universe. As a consequence, the density profile asymptotically approaches the NFW profile [34, 35] commonly observed in CDM.
@@
To characterize the properties of DM halos in simulations, it is common to fit the density profiles with certain functional choices based on the core configuration [e.g. 36, 37, 32, 38, 39, 40]. Efforts to derive analytical density profiles based on the isothermal-core configuration have also been made, either by enforcing the isothermal condition in the Jeans equation [e.g. 41, 42, 25] or within the more complex gravothermal fluid formalism [e.g. 32]. However, despite the advantage of directly capturing the isothermal core and being robust in accommodating baryonic configurations, these profiles lack a closed-form functional expression. Here, we introduce a novel functional form for the density profile that provides a close approximation to the isothermal-core configuration. Section II details the different contemporary density profiles, as well as the motivation and derivation of our profile. We also examine the analytical behavior of each density profile and the corresponding velocity dispersion profiles. Section III inspects the fitting and reconstructing results of the profiles in the context of isolated N-body simulations. We conclude the investigation with a summary in Section IV and discuss potential further applications of the profiles in future SIDM studies.
@@ II.1 Core density profiles
where $M_{\rm{NFW}}^{\prime}\left(r\right)$ is the NFW mass profile derived from the corresponding NFW scale density $\rho_{\rm{s}}^{\prime}$ and scale radius $r_{\rm{s}}^{\prime}$ [^2] Originally, the Read profile was used to characterize the baryon-induced flat core of DM halos, with an additional exponent $\alpha$ in the $\tanh^{\alpha}{\left(r/r_{\rm{c}}\right)}$ term, controlling the core flatness. However, it has also been used in the context of SIDM as a convenient fitting function [e.g. 25, 32, 39].. The characteristic core density $\rho_{\rm{c}}$ of the Read profile can be recovered as $\rho_{\rm{c}}=1.5\rho_{\rm{s}}^{\prime}r_{\rm{s}}^{\prime}/r_{\rm{c}}$. It must be noted here that the scale density and radius of the Read and other profiles (indicated by the prime notation to be distinguished from the NFW profile parameters) are not necessarily the same as the initial NFW profile parameters, with their values changing throughout the evolution of the halo.
@@ II.2 Isothermal-core density profile
To construct a density profile that satisfies the isothermal-core configuration, we start from the Jeans equation [e.g. 41, 42, 25]
$\frac{\partial}{\partial r}\Big[\rho(r)\sigma_{\rm{1D}}^{2}(r)\Big]$ $=-\rho(r)\frac{\partial\Phi(r)}{\partial r},$ (6)
where we assume spherical symmetry and isotropy. Considering only dark matter, the gravitational potential term can be written as $\partial\Phi(r)/\partial r=GM(r)/r^{2}$, with $M(r)$ as the enclosed mass profile, obtained from $\partial M(r)/\partial r=4\pi r^{2}\rho(r)$. [^3] More complicated systems involving both baryonic and dark matter would require treatments utilizing Poisson’s equation. The one-dimensional velocity dispersion $\sigma_{\rm{1D}}$ is taken to be constant within a specified radius, beyond which the profile transitions to an NFW form. This method is relatively accurate and can accommodate adiabatic contractions caused by baryons. However, no closed-form expression exists for the density profile calculated from this approach, as it necessitates numerical calculations and a more complicated boundary-stitching procedure [25].
As our goal is to arrive at an analytical density profile that closely satisfies the isothermal-core configuration while maintaining a well-defined functional form, we take a similar but slightly more simplistic approach. First and foremost, we require our density profile to exhibit similar behaviors to the profiles detailed in Section II.1, that is: having a radius-independent density at lower radii, a transition index of $n$ in the intermediate region, and a power-law of $r^{-3}$ similar to the NFW profile at the halo edge. In agreement with simulation results, we assume a constant core velocity dispersion $\sigma_{\rm{c}}$ within the region of $r\lesssim r_{\rm{c}}$, leading to the following dimensionless set of equations
$\frac{\partial\bar{M}(\bar{r})}{\partial\bar{r}}$ $=\hskip 2.84544pt\bar{r}^{2}\bar{\rho}(\bar{r}),$ (7)
$\frac{\partial\bar{\rho}(\bar{r})}{\partial\bar{r}}$ $=-\frac{1}{I}\frac{\bar{\rho}(\bar{r})\bar{M}(\bar{r})}{\bar{r}^{2}},$ (8)
with, $\bar{r}=r/r_{\rm{c}}$, $\bar{\rho}(\bar{r})=\rho(r)/\rho_{\rm{c}}$, and $\rho$ being our desired density profile. Here, $\bar{M}(\bar{r})=M(r)/4\pi r_{\rm{c}}^{3}\rho_{\rm{c}}$ and $I=\sigma_{\rm{c}}^{2}/4\pi G\rho_{\rm{c}}r_{\rm{c}}^{2}$. This parameter $I$ can be interpreted as a characteristic core kinetic-to-gravitational (K–G) ratio under the isothermal-core assumption, which can be calculated from $\bar{\rho}$ as
@@ II.3 Analytical characteristics of density profiles
Figure 2: The velocity dispersion profiles reconstructed from the density profiles presented in Figure 1 using the spherically symmetric Jeans equation (Equation 6). Velocity dispersions and radii are scaled with $\sqrt{G\,\rho_{\rm{c}}\,r_{\rm{c}}^{2}}$ and $r_{\rm{c}}$, respectively. Again, the Read (top left), Robertson-Fischer (top right), Yang (bottom left), and our (bottom right) profiles are shown with three choices of $r_{\rm{s}}/r_{\rm{c}}=3,10,30$ (in pink, green, and blue), as well as three choices of $n=1,1.75,2.5$ (in dotted, dashed, and solid lines) for the relevant profiles. The sub-panels present a zoomed-in view of the velocity dispersion profiles within the expected isothermal region of the core $r\lesssim 3\,r_{\rm{c}}$.
@@
Figure 2 shows the normalized one-dimensional velocity dispersion profiles $\bar{\sigma}_{\rm{1D}}=\sigma_{\rm{1D}}/\sqrt{G\bar{\rho}_{\rm{c}}\bar{r}_{\rm{c}}^{2}}$ reconstructed from Figure 1’s density profiles using the spherically symmetric Jeans equation (Equation 6). The sub-panels display the velocity dispersion profiles in linear scale, with a zoomed-in focus on the $r\lesssim 3\,r_{\rm{c}}$ region, where the profile is expected to exhibit isothermal behavior. For the Robertson-Fischer and Yang profiles, the velocity dispersion profiles with higher values of $n$ tend to approach the isothermal-core configuration in the innermost region of $r\ll r_{\rm{c}}$, albeit with relatively abrupt core transitions. In contrast, our profile exhibits smoother transitions across parameter choices, with the velocity dispersion remaining nearly constant throughout the core, indicating a closer approximation to the isothermal-core configuration.
@@ III.1 Simulation data & analysis
To directly measure the core density $\rho_{\rm{core}}$ and velocity dispersion $\sigma_{\rm{core}}$ for each snapshot, we start from the first radial bin $i$ with the cumulative particle count of $N_{i}\geq N_{\rm{min}}$. For the core density, we choose $N_{\rm{min}}=10^{3}$, while for the core velocity dispersion, $N_{\rm{min}}=10^{4}$. These choices are empirical and depend primarily on the simulation resolution. From the radial bin $i$, the cumulative average density $\bar{\rho}_{i}\pm\Delta{\bar{\rho}_{i}}$ is calculated. Comparing this to the density of the following radial bin $\rho_{i+1}\pm\Delta\rho_{i+1}$, if $\lvert\bar{\rho}_{i}-\rho_{i+1}\rvert\,\geq\left(\Delta{\bar{\rho}_{i}^{2}}+\Delta{\rho_{i+1}^{2}}\right)^{1/2}$, take $\bar{\rho}_{i}$ as the core density $\rho_{\rm{core}}$, else, consider the radial bin $i+1$ and repeat. For the core velocity dispersion $\sigma_{\rm{core}}$, we repeat the same process using the one-dimensional velocity dispersion profile $\sigma_{\rm{1D}}(r)$. For improved statistics, we also utilize a Monte Carlo (MC) approach, perturbing the density and one-dimensional velocity dispersion profiles according to their uncertainties, generating a set of samples. We then repeat the aforementioned process for the values of $\rho_{\rm{core}}$ and $\sigma_{\rm{core}}$ in each sample. The final results, derived from the distributions of sampled values, are consistent with those obtained directly from the original simulation configurations. The core half-density radius $r_{\rho/2}$ is obtained using a similar MC method.
@@ III.2 Fitting results
In the current picture, $C=\mathcal{O}\left(1\right)$ often takes the value of $C\simeq 0.75$ to match the results of fluid models [e.g. 17, 32, 46] and N-body simulations [e.g. 47, 17, 19, 48]. Here, we use the value of $C\simeq 0.85$ in order to approximate $\tau$ as the onset of the gravothermal catastrophe, i.e. when the core density and velocity dispersion rapidly diverge while the core radius (and mass) vanishes.
@@
Figure 3 shows the density profile fits and the reconstructed one-dimensional velocity dispersion profiles for three different snapshots of the main halo. The first snapshot, $T\simeq 0.02\,\tau$, corresponds to the earliest phase of core formation. The second, $T\simeq 0.45\,\tau$, represents a typical snapshot during core collapse, where self-similarity is most relevant. The results for this snapshot are representative of the halo’s behavior within the time interval of $T\simeq 0.05\text{--}0.85\,\tau$. The final snapshot, $T\simeq 0.94\,\tau$, shows the halo during the gravothermal catastrophe. The density profiles from each snapshot are fitted using the Read, Robertson–Fischer, Yang, and our profiles. Corresponding velocity dispersion profiles are then reconstructed based on the fitted parameters using the spherically symmetric Jeans equation (Equation 6). For reference, the directly measured core density $\rho_{\rm{core}}$ and velocity dispersion $\sigma_{\rm{core}}$ are also displayed. We find that our profile and the Yang profile generally provide good fits to the halo density structures, while the Read profile deviates significantly after the earliest snapshots. This becomes more apparent in the logarithmic slope presented in Figure 6. Additionally, although the Robertson–Fischer profile provides a reasonable fit during the earlier evolutionary stages, it fails to accurately capture the density structure’s behavior in the later phases of core collapse. The Yang profile and our profile yield notably similar results in fitting the density structures, with only minor deviations observed in the constant-density cores and the intermediate transition regions, i.e., the regions where density starts to decrease. However, these small deviations result in very different behaviors in the velocity dispersion profiles, most clearly seen in the middle panel of Figure 3, with only our profile closely matching simulations and maintaining consistency with the isothermal-core configuration. It should be noted, nonetheless, that even our profile exhibits slight deviations from the simulated velocity dispersion around $r_{\rm{s}}$. The underlying reason can be inferred from the deviations in the logarithmic slope, which also indicate that achieving a more accurate reconstruction of the velocity dispersion profile would require adopting a more complex functional form with additional degrees of freedom. The exceptions for the general goodness of fit observation occur only at the earliest snapshots, where the Yang and other profiles provide a better description of the velocity dispersion structures, and at the later stages of core collapse, where none of the profiles can accurately capture the isothermal core.
@@ Appendix A Analytical calculation of the core velocity dispersion via the isothermal K-G ratio
To inspect the core isothermal K-G ratio (Equation 9) of our density profile, we first separate $\bar{\rho}_{\rm{T25}}\left(r\right)$ into
$g(\bar{r})$ $=\left(\frac{\tanh{\bar{r}}}{\bar{r}}\right)^{n},$ (15)
$h(\bar{r})$ $=\frac{1}{\left(1+\left(\lambda\bar{r}\right)^{\gamma}\right)^{\frac{3-n}{\gamma}}},$ (16)
with $\lambda=r_{\rm{c}}/r_{\rm{s}}^{\prime}$. Here, $g(\bar{r})$ and $h(\bar{r})$ represent the approximate density profile of the halo inner part and the modifier for the halo outer edge. Taking the $\bar{r}\ll 1$ approximations, we obtain the first and second derivatives
$g^{\prime}(\bar{r})$ $=-\frac{2n}{3}\bar{r}+\mathcal{O}(\bar{r}^{3}),$ (17)
$g^{\prime\prime}(\bar{r})$ $=-\frac{2n}{3}+\frac{4n^{2}}{9}\bar{r}^{2}+\mathcal{O}(\bar{r}^{3}),$ (18)
$h^{\prime}(\bar{r})$ $=-\left(3-n\right)\frac{\left(\lambda\bar{r}\right)^{\gamma}}{\bar{r}}+\mathcal{O}(\bar{r}^{2\gamma-1}),$ (19)
$h^{\prime\prime}(\bar{r})$ $=-\left(3-n\right)\left(\gamma-1\right)\frac{\left(\lambda\bar{r}\right)^{\gamma}}{\bar{r}^{2}}+\mathcal{O}(\bar{r}^{2\gamma-2}).$ (20)

### v4

# Novel density profile for isothermal cores of dark matter halos

@@ I Introduction
Following the initial thermalization phase and throughout the gravothermal collapse process, until reaching the short-mean-free-path regime when heat conduction is only effective at the surface of the core, the halo core exhibits self-similarity in both the density and velocity dispersion profiles. That is, throughout its evolution, the halo profiles can be characterized by specific values of scale density, radius, and velocity dispersion. This behavior has been well described theoretically [e.g., 29] and confirmed in simulations for both velocity-independent and velocity-dependent self-interacting cross sections [e.g., 31, 32]. In addition, when these scale quantities are normalized by their initial values (or by the corresponding NFW values), and time is rescaled by a collapse timescale (inversely proportional to the self-interaction cross section), the evolution of halos with different parameters would become almost identical. Such universality, discussed and observed in [33], as well as the aforementioned self-similar natures, provide the foundation for adopting particular functional forms to describe the density structure of halos. In this self-similar regime, the halo core develops characteristic plateaus in both the density and velocity dispersion profiles. The emergence of a central plateau in the density profile, commonly referred to as the “core configuration”, is a key feature of SIDM models and is instrumental in addressing the core-cusp problem, as discussed above. In this context, we also define the “isothermal-core configuration” as a state in which both the density and velocity dispersion remain approximately constant within the central region of the halo. Such behaviors of SIDM have been found in cosmological and isolated N-body simulations [e.g. 34, 22, 23, 24, 32]. In the outer regions of the halo, however, the number density and velocity dispersion of DM particles are significantly smaller than in the core, resulting in negligible scattering interactions over the age of the Universe. As a consequence, the density profile asymptotically approaches the NFW profile [35, 36] commonly observed in CDM.
@@
To characterize the properties of DM halos in simulations, it is common to fit the density profiles with certain functional choices based on the core configuration [e.g. 37, 38, 32, 39, 40, 41]. Efforts to derive analytical density profiles based on the isothermal-core configuration have also been made, either by enforcing the isothermal condition in the Jeans equation [e.g. 42, 43, 25] or within the more complex gravothermal fluid formalism [e.g. 32]. However, despite the advantage of directly capturing the isothermal core and being robust in accommodating baryonic configurations, these profiles lack a closed-form functional expression. Here, we introduce a novel functional form for the density profile that provides a close approximation to the isothermal-core configuration. Section II details the different contemporary density profiles, as well as the motivation and derivation of our profile. We also examine the analytical behavior of each density profile and the corresponding velocity dispersion profiles. Section III inspects the fitting and reconstructing results of the profiles in the context of isolated N-body simulations. We conclude the investigation with a summary in Section IV and discuss potential further applications of the profiles in future SIDM studies.
@@ II.1 Core density profiles
where $M_{\rm{NFW}}^{\prime}\left(r\right)$ is the NFW mass profile derived from the corresponding NFW scale density $\rho_{\rm{s}}^{\prime}$ and scale radius $r_{\rm{s}}^{\prime}$ [^2] Originally, the Read profile was used to characterize the baryon-induced flat core of DM halos, with an additional exponent $\alpha$ in the $\tanh^{\alpha}{\left(r/r_{\rm{c}}\right)}$ term, controlling the core flatness. However, it has also been used in the context of SIDM as a convenient fitting function [e.g. 25, 32, 40].. The characteristic core density $\rho_{\rm{c}}$ of the Read profile can be recovered as $\rho_{\rm{c}}=1.5\rho_{\rm{s}}^{\prime}r_{\rm{s}}^{\prime}/r_{\rm{c}}$. It must be noted here that the scale density and radius of the Read and other profiles (indicated by the prime notation to be distinguished from the NFW profile parameters) are not necessarily the same as the initial NFW profile parameters, with their values changing throughout the evolution of the halo.
@@ II.2 Isothermal-core density profile
To construct a density profile that satisfies the isothermal-core configuration, we start from the Jeans equation [e.g. 42, 43, 25] {align} ∂∂r [ρ(r) σ_1D^2(r)] = - ρ(r) ∂Φ(r)∂r, where we assume spherical symmetry and isotropy. Considering only dark matter, the gravitational potential term can be written as $\partial\Phi(r)/\partial r=GM(r)/r^{2}$, with $M(r)$ as the enclosed mass profile, obtained from $\partial M(r)/\partial r=4\pi r^{2}\rho(r)$. [^3] More complicated systems involving both baryonic and dark matter would require treatments utilizing Poisson’s equation. The one-dimensional velocity dispersion $\sigma_{\rm{1D}}$ is taken to be constant within a specified radius, beyond which the profile transitions to an NFW form. This method is relatively accurate and can accommodate adiabatic contractions caused by baryons. However, no closed-form expression exists for the density profile calculated from this approach, as it necessitates numerical calculations and a more complicated boundary-stitching procedure [25].
As our goal is to arrive at an analytical density profile that closely satisfies the isothermal-core configuration while maintaining a well-defined functional form, we take a similar but slightly more simplistic approach. First and foremost, we require our density profile to exhibit similar behaviors to the profiles detailed in Section II.1, that is: having a radius-independent density at lower radii, a transition index of $n$ in the intermediate region, and a power-law of $r^{-3}$ similar to the NFW profile at the halo edge. In agreement with simulation results, we assume a constant core velocity dispersion $\sigma_{\rm{c}}$ within the region of $r\lesssim r_{\rm{c}}$, leading to the following dimensionless set of equations {align} ∂¯M(¯r)∂¯r =   ¯r^2 ¯ρ (¯r),
∂¯ρ(¯r)∂¯r = - 1I ¯ρ(¯r) ¯M(¯r)¯r2, with, $\bar{r}=r/r_{\rm{c}}$, $\bar{\rho}(\bar{r})=\rho(r)/\rho_{\rm{c}}$, and $\rho$ being our desired density profile. Here, $\bar{M}(\bar{r})=M(r)/4\pi r_{\rm{c}}^{3}\rho_{\rm{c}}$ and $I=\sigma_{\rm{c}}^{2}/4\pi G\rho_{\rm{c}}r_{\rm{c}}^{2}$. This parameter $I$ can be interpreted as a characteristic core kinetic-to-gravitational (K–G) ratio under the isothermal-core assumption, which can be calculated from $\bar{\rho}$ as
@@ II.3 Analytical characteristics of density profiles
Figure 2: The velocity dispersion profiles reconstructed from the density profiles presented in Figure 1 using the spherically symmetric Jeans equation (Equation II.2). Velocity dispersions and radii are scaled with $\sqrt{G\,\rho_{\rm{c}}\,r_{\rm{c}}^{2}}$ and $r_{\rm{c}}$, respectively. Again, the Read (top left), Robertson-Fischer (top right), Yang (bottom left), and our (bottom right) profiles are shown with three choices of $r_{\rm{s}}/r_{\rm{c}}=3,10,30$ (in pink, green, and blue), as well as three choices of $n=1,1.75,2.5$ (in dotted, dashed, and solid lines) for the relevant profiles. The sub-panels present a zoomed-in view of the velocity dispersion profiles within the expected isothermal region of the core $r\lesssim 3\,r_{\rm{c}}$.
@@
Figure 2 shows the normalized one-dimensional velocity dispersion profiles $\bar{\sigma}_{\rm{1D}}=\sigma_{\rm{1D}}/\sqrt{G\bar{\rho}_{\rm{c}}\bar{r}_{\rm{c}}^{2}}$ reconstructed from Figure 1’s density profiles using the spherically symmetric Jeans equation (Equation II.2). The sub-panels display the velocity dispersion profiles in linear scale, with a zoomed-in focus on the $r\lesssim 3\,r_{\rm{c}}$ region, where the profile is expected to exhibit isothermal behavior. For the Robertson-Fischer and Yang profiles, the velocity dispersion profiles with higher values of $n$ tend to approach the isothermal-core configuration in the innermost region of $r\ll r_{\rm{c}}$, albeit with relatively abrupt core transitions. In contrast, our profile exhibits smoother transitions across parameter choices, with the velocity dispersion remaining nearly constant throughout the core, indicating a closer approximation to the isothermal-core configuration.
@@ III.1 Simulation data & analysis
[…]
@@ III.2 Fitting results
In the current picture, $C=\mathcal{O}\left(1\right)$ often takes the value of $C\simeq 0.75$ to match the results of fluid models [e.g. 17, 32, 33] and N-body simulations [e.g. 47, 17, 19, 48]. Here, we use the value of $C\simeq 0.85$ in order to approximate $\tau$ as the onset of the gravothermal catastrophe, i.e. when the core density and velocity dispersion rapidly diverge while the core radius (and mass) vanishes.
@@
Figure 3 shows the density profile fits and the reconstructed one-dimensional velocity dispersion profiles for three different snapshots of the main halo. The first snapshot, $T\simeq 0.02\,\tau$, corresponds to the earliest phase of core formation. The second, $T\simeq 0.45\,\tau$, represents a typical snapshot during core collapse, where self-similarity is most relevant. The results for this snapshot are representative of the halo’s behavior within the time interval of $T\simeq 0.05\text{--}0.85\,\tau$. The final snapshot, $T\simeq 0.94\,\tau$, shows the halo during the gravothermal catastrophe. The density profiles from each snapshot are fitted using the Read, Robertson–Fischer, Yang, and our profiles. Corresponding velocity dispersion profiles are then reconstructed based on the fitted parameters using the spherically symmetric Jeans equation (Equation II.2). For reference, the directly measured core density $\rho_{\rm{core}}$ and velocity dispersion $\sigma_{\rm{core}}$ are also displayed. We find that our profile and the Yang profile generally provide good fits to the halo density structures, while the Read profile deviates significantly after the earliest snapshots. This becomes more apparent in the logarithmic slope presented in Figure 6. Additionally, although the Robertson–Fischer profile provides a reasonable fit during the earlier evolutionary stages, it fails to accurately capture the density structure’s behavior in the later phases of core collapse. The Yang profile and our profile yield notably similar results in fitting the density structures, with only minor deviations observed in the constant-density cores and the intermediate transition regions, i.e., the regions where density starts to decrease. However, these small deviations result in very different behaviors in the velocity dispersion profiles, most clearly seen in the middle panel of Figure 3, with only our profile closely matching simulations and maintaining consistency with the isothermal-core configuration. It should be noted, nonetheless, that even our profile exhibits slight deviations from the simulated velocity dispersion around $r_{\rm{s}}$. The underlying reason can be inferred from the deviations in the logarithmic slope, which also indicate that achieving a more accurate reconstruction of the velocity dispersion profile would require adopting a more complex functional form with additional degrees of freedom. The exceptions for the general goodness of fit observation occur only at the earliest snapshots, where the Yang and other profiles provide a better description of the velocity dispersion structures, and at the later stages of core collapse, where none of the profiles can accurately capture the isothermal core.
@@ Appendix A Analytical calculation of the core velocity dispersion via the isothermal K-G ratio
To inspect the core isothermal K-G ratio (Equation 6) of our density profile, we first separate $\bar{\rho}_{\rm{T25}}\left(r\right)$ into {align} g(¯r) = (tanh¯r¯r)^n,
h(¯r) = 1(1 + (λ¯r)γ)3-nγ, with $\lambda=r_{\rm{c}}/r_{\rm{s}}^{\prime}$. Here, $g(\bar{r})$ and $h(\bar{r})$ represent the approximate density profile of the halo inner part and the modifier for the halo outer edge. Taking the $\bar{r}\ll 1$ approximations, we obtain the first and second derivatives {align} g^′ (¯r) = - 2 n3 ¯r + O(¯r^3),
g^′′ (¯r) = - 2 n3 + 4n29 ¯r^2 + O(¯r^3),
h^′ (¯r) = - (3-n) (λ¯r)γ¯r + O(¯r^2γ-1),
h^′′ (¯r) = - (3-n) (γ- 1) (λ¯r)γ¯r2 + O(¯r^2γ-2).